OpenStudy (anonymous):

For the function f(x)= x^2 + 2 Find the average rate of change over the interval [1,5]

OpenStudy (anonymous):

f(5) - f(1 /(5-1) = 27-3 / 4 = 24/4 =6

OpenStudy (anonymous):

so the equation is f(y) - f(x/(f(x))